17. Foscarnet Sodium (Foscavir) 90 mg/kg is ordered q 12 hours. After dilution with D5W, the concentration yields a 12 mg/ml solution. The patient weighs 118 lIbs. Pharmacy labels the Foscarnet to be infused in 2 hours. a. How many mg will be administered per dose? Answer b. What is the volume of medication to administer per dose? Answer c. Use the answer above (in thousandth) to calculate the flow rate in ml/hr. Answer
Answer #1

Given weight: 118lb= 53.5 kg

Frequency: 12 hourly

Ordered 90 mg/kg

53.5 kg=?

53.5*90/1=4815mg (4.8 gm)

a)4815 mg of Foscavir will be administered

b)Given 12 mg/ml

Then 4815 mg=??

4815*1/12 =401.25 ml will be administered per dose

c)Flow rate: Total volume*15/Total hour*60

=401.25*15/12*60

=8.3 that is 8 drops /minute

Now

8=401.25*15/1*60

=12.5 ml/hour